A biologist has a 30% salt solution and a 60% salt soution. How many gallons of the 30% salt solution should be added to the 60% salt solution to make 24 gallons of a 40% salt solution?
Answer by stanbon(75887)

A biologist has a 30% salt solution and a 60% salt soution. How many gallons of the 30% salt solution should be added to the 60% salt solution to make 24 gallons of a 40% salt solution?
---------------
Equation:
salt + salt = salt
0.30x + 0.60(24-x) = 0.40*24
-----------------------
30x + 60*24 - 60x = 40*24
-30x = -20*24
x = (2/3)24
x = 16 gallons (amt. of 30% solution needed)
24-x = 8 gallons (amt. of 60% solution needed)
-----------
Cheers,
Stan H.
------------